# Nonlinear Transport Signatures of Hidden Symmetry Breaking in a Weyl Altermagnet

### v1

@@ Abstract
$\partial_{i}\equiv\partial_{k_{i}}$, $\mathcal{A}_{l}^{nm}=\langle u_{n}|i\partial_{l}|u_{m}\rangle$ is the Berry connection, $\Omega^{li}_{n}$ is the Berry curvature, $G^{jl}_{n}$ is the band-normalized quantum metric, and $f_{n}$ is the Fermi-Dirac function regarding the band $\varepsilon_{n}$. The relaxation time $\tau_{0}=58$ fs is adopted for Ca3Ru2O7 from experiment [40]. Because quantum metric has an extra energy denominator in Eq. 5, the quantum metric can be much more efficiently enhanced by the gapless or small-gap regions than the Berry curvature (Eq. 4).
Three terms in Eq. 3 correspond to nonlinear Drude, BCD, and QMD, respectively, which can be classified into $\mathcal{T}$-even and $\mathcal{T}$-odd parts. The $\mathcal{T}$-odd contribution (Drude and QMD) exists only when $\mathcal{T}$ or $\tau\mathcal{T}$ symmetry is broken, generating both NLAHE and nonlinear longitudinal conductivity (see Fig. 4a,b). By contrast, the $\mathcal{T}$-even contribution–BCD–appears in general but only generates NLAHE. Therefore, the $Bb2_{1}m$ phase can only exhibit NLAHE that corresponds to $\sigma_{y;xx}$ and $\sigma_{y;zz}$ according to the magnetic space group symmetry. Here, an electric field is applied inside the glide plane ($\mathcal{G}_{x}$ or $\mathcal{G}_{z}$) and the induced nonlinear current emerges along the polar axis ($\hat{y}$). But $Pn2_{1}a$ can present both NLAHE ($\sigma_{y;xx}$ and $\sigma_{y;zz}$) and nonlinear conductivity along the polar axis ($\sigma_{y;yy}$).
@@
We build the material-specific Wannier tight-binding Hamiltonian extracted from DFT calculations and then calculate nonlinear conductivity following Eq. 3. As shown in Fig. 4c-f, Weyl rings significantly enhance BCD and QMD in the band structure. In the $Bb2_{1}m$ phase, the $\{\mathcal{T}|\tau\}$ symmetry enforces a symmetric distribution of BCD but an antisymmetric distribution of QMD (Fig. 4c, 4d). Thus, the total QMD contribution ($\sigma_{y;yy}$) vanishes. We stress that near Weyl rings, QMD-induced conductivity surpasses BCD contribution by several orders of magnitude because the quantum metric is more efficiently enhanced by Weyl rings than Berry curvature. In the $Pn2_{1}a$ phase, a weak perturbation that breaks the QMD antisymmetry (see Fig. 4f) can generate a giant net QMD, leading to around one thousand times larger nonlinear conductivities than in the $Bb2_{1}m$ phase (Fig. 4g-h).
@@
The Weyl chain empowered nonlinear conductivities can be further demonstrated from their chemical potential dependence in Figs. 4g, 4h. In the energy range of Weyl chains, the nonlinear conductivities are generally strong with characteristic peaks. Compared to Weyl points, Weyl chains present more hot spots of BCD and QMD simply because Weyl rings span much larger spaces than Weyl points in the Brillouin zone. Besides Weyl chains, generic small-gap regions in the band structure also produce large QMD. In addition, the BCD peak near the Fermi energy in the $Bb2_{1}m$ phase is caused by another band crossing (see Supplementary Fig. S2). In the $Pn2_{1}a$ phase, the nonlinear conductivities are dominated by QMD compared to BCD and the Drude term (Supplementary Fig. S3).
We predict that the transition from $Bb2_{1}m$ to $Pn2_{1}a$ comes with a significant enhancement of the nonlinear transport, providing a sensitive probe to the structural distortion and resultant band structure change. The existence of longitudinal nonlinear response will be a definitive proof for the $Pn2_{1}a$ phase.
@@
Summary.—We have predicted that Ca3Ru2O7 is an AFM Weyl chain semimetal that hosts topological surface states and nonlinear transport. Compared to the $Bb2_{1}m$ phase, the hidden symmetry breaking in the $Pn2_{1}a$ phase drives the system into an altermagnetic state, deforms the Weyl chain structure and associated surface states, and generates a unique longitudinal nonlinear transport signal and significantly enhanced NLAHE. Altermagnets are usually characterized by the nonrelativistic split and anomalous Hall effect in experiments, as reported in centrosymmetric systems like CrSb [41, 42]. The $Pn2_{1}a$ phase of Ca3Ru2O7 belongs to a family of inverison-breaking altermagnet [43, 44]. We point out that the QMD-induced nonlinear response provides a sensitive diagnostic tool for identifying the altermagnet in polar systems, which complements other detection methods. More broadly, our work establishes a connection between crystal geometry and quantum geometry in correlated systems. The sensitivity of nonlinear transport to hidden orders presents new opportunities to explore unknown symmetry-breaking as a new metrology beyond the limitations of conventional crystallography or spectroscopy techniques.
@@
Note added.— A very recent experiment has measured the longitudinal and transverse nonlinear transport, as we predicted, and confirmed the $Pn2_{1}a$ phase as the ground state structure for Ca3Ru2O7 below 30 K [40].
[…]
The nonlinear transport conductivities are calculated by integrating the quantum geometry on the Fermi surface (see Supplementary Section A). In the $Bb2_{1}m$ phase, due to the presence of $\tau\mathcal{T}$, $\mathcal{G}_{x}$ and $\mathcal{G}_{z}$ symmetries, we only need to integrate the one-eighth of the Brillouin zone with a $k$-point mesh of $400\times 400\times 200$. In the $Pn2_{1}a$ phase, $\mathcal{G}_{x}$ and $\mathcal{G}_{z}$ symmetries divide the Brillouin zone into four parts, so we need to integrate a quarter of the Brillouin zone with a $k$-point mesh of $400\times 800\times 200$ ($k_{x},k_{z}\in[0,0.5],k_{y}\in[-0.5,0.5]$).
@@
Acknowledgments B.Y. and Z.Q.M. acknowledge the National Science Foundation through the Penn State Materials Research Science and Engineering Center (MRSEC) DMR 2011839. B.Y. acknowledges the financial support by the Israel Science Foundation (ISF No. 2932/21, No. 2974/23).

### v2

@@ III.5 Nonlinear transport
$\partial_{i}\equiv\partial_{k_{i}}$, $\mathcal{A}_{l}^{nm}=\langle u_{n}|i\partial_{l}|u_{m}\rangle$ is the Berry connection, $\Omega^{li}_{n}$ is the Berry curvature, $G^{jl}_{n}$ is the band-normalized quantum metric, and $f_{n}$ is the Fermi-Dirac function regarding the band $\varepsilon_{n}$. The relaxation time $\tau_{0}=58$ fs is adopted by extracting the carrier mobilities and densities from the magnetotransport experiment [47]. Because the quantum metric has an extra energy denominator in Eq. 4, the quantum metric can be much more efficiently enhanced by the gapless or small-gap regions than the Berry curvature (Eq. 3).
Three terms in Eq. 2 correspond to nonlinear Drude, BCD, and QMD, respectively, which can be classified into $\mathcal{T}$-even and $\mathcal{T}$-odd parts. The $\mathcal{T}$-odd contribution (Drude and QMD) exists only when $\mathcal{T}$ or $\tau\mathcal{T}$ symmetry is broken, generating both NLAHE and nonlinear longitudinal conductivity (see Fig. 4a, 4b). By contrast, the $\mathcal{T}$-even contribution–BCD–appears in general but only generates NLAHE. Therefore, the $Bb2_{1}m$ phase can only exhibit NLAHE that corresponds to $\sigma_{y;xx}$ and $\sigma_{y;zz}$ according to the magnetic space group symmetry. Here, an electric field is applied inside the glide plane ($\mathcal{G}_{x}$ or $\mathcal{G}_{z}$) and the induced nonlinear current emerges along the polar axis ($\hat{y}$). But $Pn2_{1}a$ can present both NLAHE ($\sigma_{y;xx}$ and $\sigma_{y;zz}$) and nonlinear conductivity along the polar axis ($\sigma_{y;yy}$).
@@
We build the material-specific Wannier tight-binding Hamiltonian extracted from DFT calculations and then calculate nonlinear conductivity following Eq. 2. As shown in Figs. 4c-4f, Weyl rings significantly enhance BCD and QMD in the band structure. In the $Bb2_{1}m$ phase, the $\{\mathcal{T}|\tau\}$ symmetry enforces a symmetric distribution of BCD but an antisymmetric distribution of QMD (Fig. 4c, 4d). Thus, the total QMD contribution ($\sigma_{y;yy}$) vanishes. We stress that near Weyl rings, QMD-induced conductivity surpasses BCD contribution by several orders of magnitude because the quantum metric is more efficiently enhanced by Weyl rings than Berry curvature. In the $Pn2_{1}a$ phase, a weak perturbation that breaks the QMD antisymmetry (see Fig. 4f) can generate a giant net QMD, leading to around one thousand times larger nonlinear conductivities than in the $Bb2_{1}m$ phase (Fig. 4g-4h).
@@
The Weyl chain empowered nonlinear conductivities can be further demonstrated from their chemical potential dependence in Figs. 4g, 4h. In the energy range of Weyl chains, the nonlinear conductivities are generally strong with characteristic peaks. Compared to an isolated Weyl point, a Weyl ring or a Weyl chain presents more hot spots of quantum geometry and leads to stronger nonlinear conductivities because it spans larger spaces than a Weyl point in the Brillouin zone and distributes in a broader energy window. Besides Weyl chains, generic small-gap regions in the band structure also produce large QMD. In addition, the BCD peak near the Fermi energy in the $Bb2_{1}m$ phase is caused by another band crossing (see SI Fig. S2 [46]). In the $Pn2_{1}a$ phase, the nonlinear conductivities are dominated by QMD compared to BCD and the Drude term (SI Fig. S3 [46]).
We propose two sensitive probes to identify the subtle structural distortion and resultant band structure change.. First, the transition from $Bb2_{1}m$ to $Pn2_{1}a$ is accompanied by a significant enhancement of the nonlinear transport, a direct manifestation of the quantum metric. This can be verified by temperature-dependent transport signatures around 48 K. Second, the emergence of a second-order longitudinal conductivity along the polar axis is definitive proof for the $Pn2_{1}a$ phase, as this response is strictly forbidden in the $Bb2_{1}m$ phase.
IV Summary
@@
We have predicted that Ca3Ru2O7 is an AFM Weyl chain semimetal that hosts topological surface states and nonlinear transport. Compared to the $Bb2_{1}m$ phase, the hidden symmetry breaking in the $Pn2_{1}a$ phase drives the system into an altermagnetic state, deforms the Weyl chain structure and associated surface states, and generates a unique longitudinal nonlinear transport signal and significantly enhanced NLAHE. Altermagnets are usually characterized by the nonrelativistic split and anomalous Hall effect in experiments, as reported in centrosymmetric systems such as CrSb [48, 49]. The $Pn2_{1}a$ phase of Ca3Ru2O7 belongs to a family of inversion-breaking altermagnet [50, 51]. We point out that the QMD-induced nonlinear response provides a sensitive diagnostic tool for identifying the altermagnet in polar systems, which complements other detection methods. More broadly, our work establishes a connection between crystal geometry and quantum geometry in correlated systems. The sensitivity of nonlinear transport to hidden orders presents new opportunities to explore unknown symmetry-breaking as a new metrology beyond the limitations of conventional crystallography or spectroscopy techniques. It is worth noting that this method can also apply to other kinds of magnets [52] (e.g., non-collinear antiferromagnets, traditional antiferromagnets, and ferromagnets).
@@
Note added. A very recent experiment has measured the longitudinal and transverse nonlinear transport, as we predicted, and confirmed the $Pn2_{1}a$ phase as the ground state structure for Ca3Ru2O7 below 30 K [47].
Acknowledgements.
@@
B.Y. and Z.Q.M. acknowledge the National Science Foundation through the Penn State Materials Research Science and Engineering Center (MRSEC) DMR 2011839. B.Y. acknowledges the financial support by the Israel Science Foundation (ISF No. 2932/21, No. 2974/23).